OpenStudy (anonymous):

write 98 as product of its prime factors

Parth (parthkohli):

\(\Large \color{purple}{ 98 = 2 * 7 * 7 } \)

OpenStudy (mertsj):

It is even, so start by dividing by 2

Parth (parthkohli):

2 * 7 = 14 14 * 7 = 98 :D

Parth (parthkohli):

@Mertsj exactly what I did :)

OpenStudy (anonymous):

For a general factorization problem, check divisibility starting from the first prime and then going down the prime list. 2,3,5,7,11, ...

OpenStudy (anonymous):

so the answer is 98=2*7*7

OpenStudy (anonymous):

more succintly written as \(2\times 7^2\)

OpenStudy (anonymous):

*succinctly